# MOFA-Video: Controllable Image Animation via Generative Motion Field Adaptions in Frozen Image-to-Video Diffusion Model

### v1

@@ 4.1 Comparison with Other State-of-the-Art Methods
4.1.1 Trajectory-based Image Animation.
For pure trajectory-based image animation, we compare our approach with DragNUWA [45]. Their method integrates the sparse motion vector into the diffusion model without employing any explicit flow and warping. Consequently, this makes it entirely uninterpretable and potentially causing various limitations. As depicted in the right portion of Fig. 8, DragNUWA incorrectly extends the influence of the user annotation to the background, causing the beach to move to the right even though the user’s intention was to control the ship’s cruise direction. This issue is particularly prone to occur in DragNUWA when the motion prior to specific parts is overly strong. In comparison, our method avoids this problem by explicitly determining the moving region through dense optical flow, thus yielding superior results. The quantitative results are shown in Tab. 2, where the proposed method shows much better performance in terms of all metrics.
@@
4.1.2 Portrait Image Animation.
We evaluate our keypoint-based MOFA-Adapter in comparison to prior state-of-the-art techniques such as StyleHEAT [44], SadTalker [50]. The qualitative results are shown in Fig. 20. As StyleHEAT [44] requires aligned images as input, we first crop and align the input image, then display the resulting output in the aligned format. We can see that StyleHEAT [44] tends to generate overly smooth outcomes and struggles with identity preservation issues due to the inherent limitations of latent inversion via StyleGAN [18, 19]. SadTalker [50] manages to process unaligned images using the paste-back technique, but it creates noticeable unnatural boundary artifacts. Additionally, it only controls head movement, leaving the target’s body static during head motion, and may produce low-resolution outputs ($256\times 256$), leading to a resolution discrepancy between facial and other parts. In comparison, our method effectively adapts to unaligned portrait images, removing noticeable paste-back artifacts, synchronizing body movements with head motion, and enabling high-resolution synthesis (up to $1024\times 1024$) results. The corresponding quantitative results are reported in Tab. 2. Our method achieves the best results in all metrics, demonstrating its superiority over existing methods.
4.1.3 User Studies.
We also conduct user studies to evaluate the effectiveness of the proposed method compared with baseline methods. For trajectory-based animation, we compare with DragNUWA using the same trajectories. For facial animation, we produce the same landmarks as SadTalker [50] using the same audio. For each task, we choose 20 samples and invite 10 volunteers to evaluate the results. For trajectory-based animation, we consider the visual qualities and the control precision in the score range of 1 to 5 (higher the better). For facial animation, we consider the fidelity to the input image, the naturalness of the motion, and the visual qualities. The results are reported in Tab. 2 and Tab. 2. We can see that the volunteers prefer the proposed method in terms of all the aspects.
@@ 4.2 Ablation Studies
4.2.1 Network Structure of MOFA-Adapter.
To evaluate the effectiveness of our network design, we compare the results of our method with three different variants:
@@
The corresponding results are in Fig. 10. It is evident that the sparse-conditioning model is unable to precisely manipulate the target object’s trajectory, as the dog does not move to the right, and the background slightly shifts to the left. We argue that this issue arises from the spatial misalignment caused by not employing the spatial warping operation. While the sparse-warping model can successfully manage the target’s trajectory, it struggles to effectively transform the deep feature from the initial frame due to the lack of dense feature control provided by dense optical flow, resulting in significant artifacts in the output. The non-tuning model, however, experiences over-control from the dense optical flow and is unable to generate suitable content using the strong prior of the SVD. In comparison, our comprehensive model achieves optimal results by integrating the Sparse-to-Dense approach, spatial feature-space warping, and domain-specific tuning.
4.2.2 Longer Video Inference.
Additionally, we evaluate our periodic sampling approach against two alternative strategies:
@@
4.2.3 The effectiveness of Domain-Aware MOFA-Adapter.
As introduced in Sec. 3, we train different MOFA-Adapters for different motion domains. To evaluate the effectiveness and the necessity of domain-specific tuning, we directly use our trajectory-based model to perform landmark-based portrait image animation. As shown in Fig. 12, directly applying the model trained on open-domain will cause unnatural facial expressions.
@@ 5 Conclusion
In this paper, we introduce a novel pipeline for generic and controllable image-to-video animation from multiple motion domains (e.g., handcrafted trajectories, dense flows, human key-points). To achieve this goal, we design the sparse-to-dense MOFA-Adapter to control the generated motions in the video generation pipeline. Powered by the proposed framework, we can achieve controllable video generation via fine-grained control which is unified as flow motion fields. The experiments show the advantage of the proposed framework over current state-of-the-art methods for various applications.
@@ 0.A.2 More Training Details
The trajectory-based model is trained on the WebVid-10M [5] dataset using the AdamW optimizer with a learning rate of $2\times 10^{-5}$. The batch size is set at 8, and the total number of training iterations is 100,000. The portrait-based model is trained on a self-compiled dataset which includes 5,889 different human portrait videos. The AdamW optimizer is used, with a learning rate of $2\times 10^{-5}$. The batch size is set to 1, and the total training iteration is 200,000.
@@ 0.A.3 Inference via Multiple MOFA-Adapters
As indicated in the main paper, we can integrate multiple MOFA-Adapters for more sophisticated and complex control using control signals from various modalities. For instance, users can merge the landmark signal with handcrafted trajectories. Specifically, we first route the trajectory control signals and landmark signals through the MOFA-Adapter for each modality separately. Contrary to the original Multi-ControlNet [48] algorithm, we employ a mask-aware strategy to define the control area for each MOFA-Adapter. Specifically, the user can designate the region where the landmark signal is accountable for, such as the human face region. Based on this mask, we extract the deep feature of the multi-scale output of the landmark-based MOFA-Adapters within the mask region, and that of the trajectory-based MOFA-Adapters outside the mask region. Finally, we input the combined features into the frozen SVD to obtain the final output.
@@ 0.C.1.1 Motion Brush
As stated in the main paper, we can employ motion mask brushes to attain detailed control by designating the spatial region of the flow patterns since our method utilizes intermediate optical flow patterns for motion control. Gen-2 [1] also supports motion brushes, but it only supports basic directions (Mask + direction) and is incapable of executing non-linear complex controls (for instance, blinking). Our method combines regional mask and trajectory (mask + trajectory), being able to handle advanced non-linear motions. The comparative results corresponding to this are displayed in Fig. 19.
@@ 0.C.2 Portrait Image Animation
More visual comparison results with StyleHEAT [44] and SadTalker [50] are demonstrated in Fig. 20. We also give more quantitative results with our methods and SadTalker [50] on visual quality (LPIPS). The proposed method shows a much better performance on visual quality (0.2099) than SadTalker (0.2308). Besides, our method also shows comparable results on lip synchronization from the same landmark generated from audio using SadTalker [50]. We give some examples in the supplementary video.
@@ Appendix 0.D Ablation Study Results
We also consider the quantitative comparison of the ablation studies in network structure. The same dataset is used for evaluation as the one used for quantitative comparisons with DragNUWA [45] in the main paper. As shown in Tab. 3, the proposed full method achieves the most balanced results in terms of all metrics. Our method w/o tuning of the MOFA-Adapter shows very limited motions compared with our full methods. Our method w/o S2D shows second-best results. However, from LPIPS, the generated video is different from the motion guidance. Finally, our method uses explicit warping as motion control, removing the explicit motion warping shows much worse results in all metrics. For video results, please refer to the video demo provided in the supplementary. We also provide the video ablation results for longer video generation and domain-aware MOFA-Adapter in the video demo.

### v3

@@ 4.1 Comparison with Other State-of-the-Art Methods
Trajectory-based Image Animation. For pure trajectory-based image animation, we compare our approach with DragNUWA [45]. Their method integrates the sparse motion vector into the diffusion model without employing explicit flow and warping. Consequently, this makes it entirely uninterpretable and potentially causing various limitations. As depicted in the right part of Fig. 8, DragNUWA incorrectly extends the influence of the user annotation to the background, causing the beach to move to the right even though the user’s intention was to control the ship’s cruise direction. This issue is particularly prone to occur in DragNUWA when the motion prior to specific parts is overly strong. In comparison, our method avoids this problem by explicitly determining the moving region through dense optical flow, thus yielding superior results. The quantitative results are shown on Tab. 2, where the proposed method shows much better performance in terms of all metrics.
@@
Portrait Image Animation. We evaluate our keypoint-based MOFA-Adapter in comparison to prior state-of-the-art techniques such as StyleHEAT [44], Sad-Talker [50]. The qualitative results are shown in Fig. 9. As StyleHEAT [44] requires aligned images as input, we first crop and align the input image, then display the resulting output in the aligned format. We can see that StyleHEAT [44] tends to generate overly smooth outcomes and struggles with identity preservation issues due to the inherent limitations of latent inversion via StyleGAN [18, 19]. SadTalker [50] manages to process unaligned images using the paste-back technique, but it creates noticeable unnatural boundary artifacts. Additionally, it only controls head movement, leaving the target’s body static during head motion, and may produce low-resolution outputs ($256\times 256$), leading to a resolution discrepancy between facial and other parts. In comparison, our method effectively adapts to unaligned portrait images, removing noticeable paste-back artifacts, synchronizing body movements with head motion, and enabling high-resolution synthesis (up to $1024\times 1024$) results. The corresponding quantitative results are reported in Tab. 2. Our method achieves the best results on all metrics, demonstrating its superiority over existing methods.
User Studies. We also conduct user studies to evaluate the effectiveness of the proposed method compared with baseline methods. For trajectory-based animation, we compare with DragNUWA using the same trajectories. For facial animation, we produce the same landmarks as SadTalker [50] using the same audio. For each task, we choose 20 samples and invite 10 volunteers to evaluate the results. For trajectory-based animation, we consider the visual qualities and the control precision in the score range of 1 to 5 (higher the better). For facial animation, we consider the fidelity to the input image, the naturalness of the motion, and the visual qualities. The results are reported in Tab. 2 and Tab. 2. We can see that the volunteers prefer the proposed method in terms of all the aspects.
@@ 4.2 Ablation Studies
Network Structure of MOFA-Adapter. To evaluate the effectiveness of our network design, we compare the results of our method with three different variants:
@@
The corresponding results are in Fig. 10. The sparse-conditioning model is unable to precisely manipulate the target object’s trajectory, as the dog does not move to the right, and the background slightly shifts to the left. We argue that this issue arises from the spatial misalignment caused by not employing the spatial warping operation. While the sparse-warping model can successfully manage the target’s trajectory, it struggles to effectively transform the deep feature from the initial frame due to the lack of dense feature control provided by dense optical flow, resulting in significant artifacts in the output. The non-tuning model, however, experiences over-control from the dense optical flow and is unable to generate suitable content using the strong prior of the SVD. In comparison, our comprehensive model achieves optimal results.
Longer Video Inference. Additionally, we evaluate our periodic sampling approach against two alternative strategies:
@@
The effectiveness of Domain-Aware MOFA-Adapter. As introduced in Sec. 3, we train different MOFA-Adapters for different motion domains. To evaluate the effectiveness and the necessity of domain-specific tuning, we directly use our trajectory-based model to perform landmark-based portrait image animation. As shown in Fig. 12, directly applying the model trained on open-domain will cause unnatural facial expressions.
@@ 5 Conclusion
In this paper, we introduce a novel pipeline for generic and controllable image-to-video animation from multiple motion domains (e.g., handcrafted trajectories, dense flows, and human key-points). To achieve this goal, we design the sparse-to-dense MOFA-Adapter to control the generated motions in the video generation pipeline. Powered by the proposed framework, we can achieve controllable video generation via fine-grained control which is unified as flow motion fields. The experiments show the advantage of the proposed framework over current state-of-the-art methods for various applications.
@@ 0.A.2 More Training Details
The trajectory-based model is trained on the WebVid-10M dataset  [5] using the AdamW optimizer with a learning rate of $2\times 10^{-5}$. The batch size is set at 8, and the total number of training iterations is 100,000. The portrait-based model is trained on a self-compiled dataset that includes 5,889 different human portrait videos. The AdamW optimizer is used, with a learning rate of $2\times 10^{-5}$. The batch size is set to 1, and the total training iteration is 200,000.
@@ 0.A.3 Inference via Multiple MOFA-Adapters
As indicated in the main paper, we can integrate multiple MOFA-Adapters for more sophisticated and complex control using control signals from various modalities. For instance, users can merge the landmark signal with handcrafted trajectories. Specifically, we first route the trajectory control signals and landmark signals through the MOFA-Adapter for each modality separately. Contrary to the original Multi-ControlNet [48] algorithm, we employ a mask-aware strategy to define the control area for each MOFA-Adapter. Specifically, the user can designate the region where the landmark signal is accountable for, such as the human face region. Based on this mask, we extract the deep feature of the multiscale output of the landmark-based MOFA-Adapters within the mask region, and that of the trajectory-based MOFA-Adapters outside the mask region. Finally, we input the combined features into the frozen SVD to obtain the final output.
@@ 0.C.1.1 Motion Brush
As stated in the main paper, we can employ motion mask brushes to attain detailed control by designating the spatial region of the flow patterns since our method utilizes intermediate optical flow patterns for motion control. Gen-2 [1] also supports motion brushes, but it only supports basic directions (Mask + direction) and is incapable of executing non-linear complex controls (for instance, blinking). Our method combines regional mask and trajectory (mask + trajectory), being able to handle advanced nonlinear motions. The comparative results corresponding to this are displayed in Fig. 19.
@@ 0.C.2 Portrait Image Animation
More visual comparison results with StyleHEAT [44] and SadTalker [50] are demonstrated in Fig. 20. We also give more quantitative results with our methods and SadTalker [50] on visual quality (LPIPS). The proposed method shows a much better performance on visual quality (0.2099) than SadTalker (0.2308). In addition, our method also shows comparable results in lip synchronization. We give some examples in the supplementary video.
@@ Appendix 0.D Ablation Study Results
We also consider the quantitative comparison of ablation studies in network structure. The same dataset is used for evaluation as the one used for quantitative comparisons with DragNUWA [45] in the main paper. As shown in Tab. 3, the proposed full method achieves the most balanced results in terms of all metrics. Our method w/o tuning of the MOFA-Adapter shows very limited motions compared with our full methods. Our method w/o S2D shows second-best results. However, from LPIPS, the generated video is different from the motion guidance. Finally, our method uses explicit warping as motion control, removing the explicit motion warping shows much worse results in all metrics. For video results, please refer to the video demo provided in the supplementary. We also provide the video ablation results for longer video generation and domain-aware MOFA-Adapter in the video demo.